The table below gives the list price and the number of bids received for five randomly selected items sold through online auctions. Using this data, consider the equation of the regression line, yˆ=b0+b1xy^=b0+b1x, for predicting the number of bids an item will receive based on the list price. Keep in mind, the correlation coefficient may or may not be statistically significant for the data given. Remember, in practice, it would not be appropriate to use the regression line to make a prediction if the correlation coefficient is not statistically significant.

Price in Dollars 2323 2626 3131 4040 4848
Number of Bids 33 44 66 77 99

Table

Copy Data

Step 1 of 6 :  

Find the estimated slope. Round your answer to three decimal places.
Answer #1

using minitab>stat>basic stat>correlation

we have

Correlation: price in doller, number of bids

Pearson correlation of price in doller and number of bids = 0.981
P-Value = 0.003

since p value is less than 0.05 so  the correlation coefficient is statistically significant.

now using minitab>stat>Regression>line plot

we have


Regression Analysis: number of bids versus price in doller

The regression equation is
number of bids = - 20.17 + 2.499 price in doller


S = 5.86798 R-Sq = 96.3% R-Sq(adj) = 95.0%


Analysis of Variance

Source DF SS MS F P
Regression 1 2655.50 2655.50 77.12 0.003
Error 3 103.30 34.43
Total 4 2758.80

the estimated slope is 2.499